EP021

PI: Patnaik
Title: B1938+666

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A/dual pol
Obs time (UT): EU 01:30(17/02)-14:00(17/02)
Stations: EbMcNtOnTrSmWb+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tugnoli Marco at Mon Feb 17 09:15:43 EST 1997) Success

Onsala (John Conway at Mon Feb 17 10:11:34 EST 1997) Minor failures
Stowed because of high winds around 10:30 UT, otherwise no known problems

Medicina (GMaccaferri at Tue Feb 18 04:58:16 EST 1997) Success
Or at least it seem to me!

Effelsberg (dgraham at Tue Feb 18 05:52:40 EST 1997) Success
No known problems known

Westerbork (Tony Foley at Tue Feb 18 10:11:05 EST 1997) Minor failures
RT8 had to bee removed from the array because of pointing problems 0315-0625UT and 0720-1346UT.

Torun (Roman Feiler at Wed Feb 19 18:32:23 EST 1997) Success
Clear sky. Only 7 BBCs used.

Jodrell (Alastair Gunn at Fri Feb 21 09:02:07 EST 1997) Minor failures
POLARISATIONS ARE CROSSED. Lost LO lock on VC06 at start of experiment, could not solve. Problem disappeared at 03:57, probably temperature related (investigating). From about 02:15 to 03:51 telescopes remained on 1842+681 due to MERLIN control problems. Tapes NGS00297,52EF JPL01225,0CE5 and JPL01550,42AD sent to Bonn via Woodcroft on 17/02/97.

Cambridge (Alastair Gunn at Fri Feb 21 09:14:02 EST 1997) Minor failures
No major problems. From about 02:15 to about 03:51 telescope remained on 1842+681 due to MERLIN control problem. Tapes EVN00024,3AAD NOAA0097,1F87 and JPL01614,66B5 sent to Bonn via Woodcroft on 17/02/97.

Calibration (Fredrik T Rantakyro at Thu May 8 11:31:47 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ep21.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Walter Alef at Wed Jan 7 12:19:12 MET 1998) Minor failures
Noto and Crimea did not observe. Wrong Jodrell polarisation was compensated at the correlator.